OpenStudy (anonymous):

oh sorry f=guys but i deleted my question by a mastake could u tell me the answer again the problem is 2y+7+-1 how do you solve this.???

OpenStudy (mertsj):

2y+7+-1=2y+6

OpenStudy (mertsj):

Add the like terms. And you are not solving it. You are simplifying.

OpenStudy (katrinakaif):

-4

OpenStudy (karatechopper):

wasnt it equal to -1? and not adding -1?

OpenStudy (anonymous):

oh yes

OpenStudy (katrinakaif):

2y + 7 = -1 (it was your original problem) 2y = -1 - 7 2y = -8 y = -4

OpenStudy (anonymous):

the problem is 2y+7=-1

OpenStudy (karatechopper):

so u want me to show u how kat got all tht?

OpenStudy (katrinakaif):

Understood?

OpenStudy (katrinakaif):

1) Your ultimate goal is to bring all the constants onto one side of the equal sign Therefore, lets move the + 7 to the other side by substracting 7 with -1, resulting -8

OpenStudy (katrinakaif):

Now your left with 2y = -8 Divide by 2 to get "y" by itself y = -8/2 y = -4 is your answer when simplified

OpenStudy (karatechopper):

nvm

OpenStudy (anonymous):

oh thx guys !! i got it now!

OpenStudy (katrinakaif):

No Problemo =)

OpenStudy (karatechopper):

gooooddd!!

OpenStudy (anonymous):

duha what grade r u in ?